# A Continuous Energy Ising Machine Leveraging Difference-of-Convex Programming

### v1

@@ 4 Methods
The authors declare that the main data supporting the findings of this study are available within the paper and its Supplementary files. Publicly available benchmark datasets used in this study, such as the G-set graphs and the Biq Mac Library, can be accessed at: https://web.stanford.edu/~yyye/yyye/Gset/ and https://biqmac.aau.at/biqmaclib.html, respectively.
5 Supplementary Materials
@@
Supplementary Figure 1: Benchmarking results for a $10^{3}$ spin Sherrington-Kirkpatrick (SK) model. The plot shows the evolution of Ising energy as a function of computation time (log scale) for several state-of-the-art solvers alongside our proposed methods. Our solvers exhibit significantly faster convergence and attain lower energy configurations compared to existing techniques. The Free Energy Machine (FEM) was evaluated using its official implementation [51] on an NVIDIA Jetson Nano.
@@
Supplementary Figure 2: Benchmark results for a $10^{4}$ spin SK model. The plot shows Ising energy versus computation time on a logarithmic scale for a range of established and proposed solvers. Our solvers demonstrate consistent advantages in both convergence speed and energy minimization as problem size increases. FEM results are obtained using the official implementation [51], executed on an NVIDIA Jetson Nano.
@@
Supplementary Figure 3: Benchmark results for a large-scale sparse Ising instance with $10^{4}$ spins and $1\%$ connectivity. Non-zero couplings $J_{ij}$ are drawn uniformly from $9$-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). Ising energy is plotted against computation time on a logarithmic scale. The results highlight the scalability and efficiency of our solvers under low-connectivity and discrete-weight regimes. All experiments were executed on a NVIDIA Jetson Nano.
@@
Supplementary Figure 4: Performance comparison on a fully connected Ising model with $10^{5}$ spins and binary couplings $J_{ij}\in\{-1,+1\}$ drawn with equal probability. Top: Ising energy evolution over wall-clock time (log scale). Bottom: Ising energy as a function of iteration count. These results demonstrate the rapid energy descent of our solvers on large and dense instances. All experiments were performed on a single NVIDIA RTX 3090 GPU.
@@
Supplementary Figure 5: Performance comparison on a fully connected $10^{5}$ spin SK. Ising energy is plotted against iteration count up to 100 steps. All methods were executed on a single NVIDIA RTX 3090 GPU.
@@
Supplementary Figure 6: Performance comparison of Ising energy for a $10^{5}$ spin Ising model with $1\%$ sparsity. Couplings $J_{ij}$ were drawn uniformly from signed $9$-bit integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). All algorithms were executed on a single NVIDIA RTX 3090 GPU.
@@
Supplementary Figure 7: Time evolution of Ising energy for a $10^{6}$ spin fully connected Ising model with structured couplings $J_{ij}=\sin(i\,j+\mathrm{seed})$. The benchmark was executed on a single NVIDIA V100 GPU.
@@
Supplementary Figure 8: Time evolution of Ising energy for a $10^{6}$ spin sparsely connected Ising model, where $0.1\%$ of the couplings are nonzero. Each active coupling $J_{ij}$ is sampled uniformly from the set of $9$-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). All solvers were run on a single NVIDIA V100 GPU.
@@
Supplementary Figure 9: Time evolution of Ising energy for a $10^{6}$ spin extremely sparse Ising model with $0.01\%$ connectivity. Each nonzero coupling $J_{ij}$ is drawn uniformly from the set of $9$-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). All solver experiments were executed on a single NVIDIA V100 GPU.
@@
Supplementary Figure 10: Benchmarking performance on a $10^{7}$ spin sparse Ising model with $0.001\%$ connectivity. Coupling coefficients $J_{ij}$ are drawn uniformly at random from the set of signed 9-bit integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). Each solver was executed using four NVIDIA V100 GPUs. The vertical axis shows the Ising energy, while the horizontal axis (log scale) denotes computation time in seconds.
@@
Supplementary Figure 11: Benchmarking performance on a $10^{8}$-spin extremely sparse Ising model with connectivity of only $0.00001\%$. Each nonzero coupling $J_{ij}$ is drawn uniformly at random from the set of signed 9-bit integers, $\{-2^{9}+1,\dots,2^{9}-1\}$. The algorithms were run on two NVIDIA H100 GPUs. The Ising energy is plotted as a function of computation time (in minutes).
@@
Supplementary Figure 12: Battery-powered edge computing setup based on the NVIDIA Jetson Nano module. The system includes a display, keyboard, and mouse interfaced with the Jetson board, powered through a $60,000$ mAh portable power system.
@@
An Ising model with $n$ spins and an external field can be equivalently represented as an $(n+1)$ spin Ising model without any external field. Consider the Ising model with an external field:
@@
It follows from (S1) and (S2) that for all $\boldsymbol{s}\in\{-1,1\}^{n}$ and $t\in\{-1,1\}$,
@@
By the definition of $\boldsymbol{\sigma}^{*}$, we have for all $\boldsymbol{s}\in\{-1,1\}^{n}$ and $t\in\{-1,1\}$,
@@
We have from (S3) and (S4) that, for any $\boldsymbol{s}\in\{-1,1\}^{n}$,
$\mathcal{E}(\boldsymbol{s}^{*})=\mathcal{E}(t_{0}\boldsymbol{s}_{0})=\hat{\mathcal{E}}(\left(\boldsymbol{s}_{0},t_{0}\right))\leqslant\hat{\mathcal{E}}(\left(\boldsymbol{s},1)\right)=\mathcal{E}(\boldsymbol{s}).$
This proves that $\boldsymbol{s}^{*}$ is a ground state of (S1). ∎
@@
To relax the discrete spin variables, we replace the binary vector $\boldsymbol{s}$ with a continuous vector $\boldsymbol{x}\in\mathbb{R}^{n}$ and define the corresponding relaxed energy:
@@
To encourage the components of $\boldsymbol{x}$ toward binary values, we introduce an attractor function:
$\mathcal{A}(\boldsymbol{x})=\frac{\beta}{4}(x_{1}^{4}+\dots+x_{n}^{4})-\frac{\alpha}{2}(x_{1}^{2}+\dots+x_{n}^{2})$ (S7)
which penalizes deviations from $\pm\sqrt{\alpha/\beta}$. The total Hamiltonian is then given by the sum
$\mathcal{H}=\mathcal{A}+\mathcal{E},$ (S8)
which forms the basis for our proposed Ising solver. The key observation behind our algorithm is that the Hamiltonian in (S8) can be rewritten as a difference of convex functions:
$\mathcal{H}=f-g,$
where
@@
The functions $f$ is convex for all $\beta>0$ and there exists $\alpha_{0}$ such that $g$ is strongly convex for all $\alpha>\alpha_{0}$.
@@
The Hamiltonian $\mathcal{H}$ is coercive for all $\beta>0$, i.e., $\mathcal{H}(\boldsymbol{x})\to+\infty$ as $\|\boldsymbol{x}\|_{2}\to\infty$.
@@
Since the Hamiltonian $\mathcal{H}$ is continuous and coercive, this follows from the Weierstrass extreme value theorem [57]. ∎
@@
$\mathcal{E}(\mathrm{sign}(\boldsymbol{x}^{*}))\leqslant\mathcal{E}(\boldsymbol{s}^{*})$ (S12)
@@
By assumption, $\boldsymbol{x}^{*}=\lambda\,\mathrm{sign}(\boldsymbol{x}^{*})$, where $\lambda>0$. Setting $\boldsymbol{x}=\lambda\boldsymbol{s}^{*}$ in (S13), we obtain
@@
Now, it follows from (S7) that $\mathcal{A}(\lambda\,\mathrm{sign}(\boldsymbol{x}^{*}))=\mathcal{A}(\lambda\boldsymbol{s}^{*})$. Additionally, we have from (S6) that $\mathcal{E}(\lambda\boldsymbol{x})=\lambda^{2}\mathcal{E}(\boldsymbol{x})$. Moreover, using (S6), we have $\mathcal{E}(\lambda\boldsymbol{x})=\lambda^{2}\mathcal{E}(\boldsymbol{x})$. Substituting these into (S14), we obtain (S12). ∎
@@
In general, the challenging aspect of convergence analysis lies in establishing that the iterates are bounded. This is often assumed [41] to facilitate the rest of the analysis. However, for DOCH, boundedness of the iterates can be established unconditionally.
@@
By Lemma 6, the sequence $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$ is monotone (nonincreasing). Hence, for all $k\geqslant 1$,
@@
As $\mathcal{H}$ is coercive, $\|\boldsymbol{x}\|\to\infty$ implies $\mathcal{H}(\boldsymbol{x})\to\infty$, so the sublevel set $\{\boldsymbol{x}:\mathcal{H}(\boldsymbol{x})\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})\}$ must be bounded. However, we know from (S21) that the entire sequence $\{\boldsymbol{x}^{(k)}\}$ belongs to this sublevel set, and hence must be bounded in $\mathbb{R}^{n}$.
@@
Since $\mathcal{H}(\boldsymbol{x}^{(k)})$ is bounded, there exists a constant $C$ such that $\mathcal{H}(\boldsymbol{x}^{(k)})\geqslant C$ for all $k\geqslant 1$. Thus, we have
$\frac{\mu}{2}\sum_{k=0}^{N}\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k+1)}\|_{2}^{2}\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})-\mathcal{H}(\boldsymbol{x}^{(N+1)})\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})-C.$
Letting $N\to\infty$, we have
$\frac{\mu}{2}\sum_{k=0}^{\infty}\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k+1)}\|_{2}^{2}\leqslant+\infty.$ (S22)
Consequently, we get (S20). ∎
We note that (S22) alone does not ensure the convergence of the sequence $\{\boldsymbol{x}^{(k)}\}$. We have to show that $\{\boldsymbol{x}^{(k)}\}$ is a Cauchy sequence [57]. We will establish this in Theorem 2 using the observation, namely that our Hamiltonian is analytic (in fact, a polynomial), and hence satisfies the Lojasiewicz gradient inequality [58]. Nevertheless, based on the analysis so far, we already have the following result.
@@
The iterates $\{\boldsymbol{x}^{(k)}\}$ are bounded by Proposition 7, and hence, by the Bolzano-Weierstrass theorem [57], have at least one limit point, say, $\boldsymbol{x}^{*}\in\mathbb{R}^{n}$. In particular, we can extract a subsequence $\{\boldsymbol{x}^{(k_{i})}\}$ such that
@@
Letting $i\to\infty$ on both sides and using (S22), we have
@@
We will now use the classical Lojasiewicz gradient inequality [58] to show that the DOCH iterates is convergent, namely that it has a unique limit point. For this, we will need the following result.
@@
Therefore, letting $\sigma=\lambda_{\max}(\mathbf{J}+\alpha\mathbf{I})>0$, we get (S27). ∎
The following is the main result concerning the convergence of the DOCH iterates. The proof is based on the analysis in [59, 60].
@@
From Lemma 6, we know that $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$ is non-increasing. Moreover, since $\boldsymbol{x}^{*}$ is a limit point of $\boldsymbol{x}^{(k)}$ and $\mathcal{H}$ is continuous,
@@
Since $\mathcal{H}$ is a real‐analytic function (in fact, a polynomial), it satisfies the Łojasiewicz gradient inequality [58, 59] in some neighbourhood of the critical point $\boldsymbol{x}^{*}$. More specifically, there exist $\theta\in[0,1),c>0$ and $r>0$, such that for all $\|\boldsymbol{x}-\boldsymbol{x}^{*}\|<r$,
@@
It will be convenient to reformulate this in terms of the function
$\varphi(t)=\frac{c}{1-\theta}\,t^{1-\theta}.$
We can verify that this function is concave and non-decreasing on $0,\infty)$, and $\varphi^{\prime}(t)=ct^{-\theta}$ for any $t>0$. In particular, we can write (S29) as
@@
We know from (S20) and (S28) that $\delta_{k}\to 0$ and $\Delta_{k}\to 0$. As $\varphi$ is non-decreasing and $\varphi(t)\to 0$ as $t\to 0$, $\varphi(\Delta_{k})\to 0$. Moreover, as $\boldsymbol{x}^{*}$ is a limit point of $\{\boldsymbol{x}^{(k)}\}$, we can find $k_{0}\geqslant 1$ such that
@@
On the other hand, we have from (S16) that
@@
Combining this with (S31), we get
@@
Therefore, we can conclude from (S31) and (S35) that
@@
That is, $\boldsymbol{x}^{k_{0}+1}$ is in the neighbourhood of $\boldsymbol{x}^{*}$ stipulated in (S29). In fact, we claim that the entire tail $\{\boldsymbol{x}^{k}\},k>k_{0}$ belongs to this neighbourhood. This can be shown using contradiction. Indded, suppose $k_{1}$ is the smallest $k>k_{0}$ such that $\|\boldsymbol{x}^{(k_{1})}-\boldsymbol{x}^{*}\|_{2}\geqslant r$. This means that the points $\boldsymbol{x}_{k_{0}},\ldots,\boldsymbol{x}_{k_{1}-1}$ are with the neighbourhood stipulated in (S29). Therefore, using (S34) repeatedly, we have for $k_{0}\leqslant k<k_{1}$,
$\delta_{k+1}\leqslant\frac{2\sigma}{\mu}\bigl{(}\varphi(\Delta_{k})-\varphi(\Delta_{k+1})\bigr{)}+(\delta_{k}-\delta_{k+1}).$
Summing both sides,

### v2

@@ 4 Methods
The authors declare that the main data supporting the findings of this study are provided within the paper and its supplementary files. The public benchmark datasets used in our experiments, including the G-set graphs and the Biq Mac Library, can be accessed at https://web.stanford.edu/~yyye/yyye/Gset/ and https://biqmac.aau.at/biqmaclib.html.
Annealing Markov Chain Monte Carlo (MCMC) SA [kirkpatrick1983, isakov2015], MARS [Yavorsky2019], NMFA [King2018] high yes poor no no hard
Simulated Hamiltonian Dynamics bSB [goto2019, goto2021] high yes good no no hard
CIM [tiunov2019, honjo2021] high yes good no no hard
SIA [jiang2024] high yes good no no hard
Continuous Optimization CP [beck2002] low no poor yes yes
SDP [Luo2010] low no poor yes yes
DOCH (present work) high yes good yes yes easy
Table 1: Comparison of key aspects of our Ising solver with existing solvers. SA: Simulated Annealing, MARS: Mean field Annealing from a Random State, NMFA: Noisy Mean Field Annealing, bSB: ballistic Bifurcation Machine, CIM: Coherent Ising Machine, SIA: Spring Ising Algorithm, CP: Convex Programming, SDP: Semidefinite Programming, DOCH: Difference of Convex Hamiltonian. (A convergence guarantee means that the iterative process will asymptotically reach a fixed point or a local minimum. An approximation guarantee, on the other hand, concerns the quality of the final solution, describing how close it is to the true ground state).
Figure 4: Evolution of Ising energy with runtime for the $10^{3}$-spin SK model. (a) Comparison of our solvers (DOCH and ADOCH) with state-of-the-art Ising solvers: Simulated Annealing (SA) [isakov2015, inagaki2016], ballistic Bifurcation Machine (bSB) [goto2021], Simulated Coherent Ising Machine (SimCIM) [goto2021], and Spring Ising Algorithm (SIA) [jiang2024]. Solid curves show the best energy values achieved across $100$ independent runs (Supplementary Notes 6 and 7 for the parameter settings). The dashed black line indicates the lowest energy obtained using the GW-SDP [goto2019, inagaki2016]. (b), (c) Histograms of $100$ Ising energies, each obtained after $1000$ iterations and with different initializations. Results obtained on a NVIDIA RTX 3050 GPU laptop.
Figure 5: Ising energy versus runtime on the $K_{2000}$ benchmark. (a) Comparison of our solvers (DOCH and ADOCH) with state-of-the-art Ising solvers. Solid curves show the mean Ising energy across $100$ independent runs (Supplementary Notes 6 and 7 for the parameter settings). The dashed black line indicates the lowest energy obtained using the GW-SDP [goto2019, inagaki2016]. (b), (c) Histograms of $100$ Ising energies, each obtained after $1000$ iterations and with different initializations. Results obtained on a NVIDIA RTX 3050 GPU laptop.
Figure 6: MAX-CUT results on Biq Mac graphs. (a) Table summarizing the graph characteristics and the average time-to-solution (Avg TTS) for each solver to reach 9999% of the maximum cut value. (Best Avg TTS values are highlighted in red, and second-best in blue). (b) Bar chart comparing the best cut values achieved after $1000$ iterations by SA [isakov2015, inagaki2016], bSB [goto2021], SimCIM [goto2021], SIA [jiang2024], and our Ising solvers DOCH and ADOCH. Results obtained on NVIDIA Jetson Nano.
Figure 7: Benchmarking solver performance at varying Ising model sizes. (a) MAX-CUT optimization on an $800$-node $G_{10}$ graph. The dashed black line marks the lowest energy found using the GW-SDP. Each solver is run for $1000$ iterations, with $100$ different initializations. The mean cut values are then plotted against the number of iterations. Results obtained on an NVIDIA Jetson Nano, with cut values plotted against iterations. (b) Benchmarking on a $10^{4}$-spin Sherrington–Kirkpatrick (SK) model, executed on NVIDIA Jetson Nano. Ising energy is shown as a function of time (log scale) in seconds. (c) Fully connected $10^{7}$-spin Ising model with couplings $J_{ij}=\sin(i\,j+\mathrm{seed})$. Experiments were conducted using four NVIDIA H100 GPUs. The energy trajectories over time are shown. (d) Sparse $10^{8}$-spin Ising model with $0.00001\%$ connectivity and 99-bit signed integer couplings ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$), executed on two NVIDIA H100 GPUs. Energy is plotted against computation time (in minutes).
5 Supplementary Material
@@
Figure S1: Benchmarking results for a $10^{3}$ spin Sherrington-Kirkpatrick (SK) model. The plot shows the evolution of Ising energy as a function of computation time (log scale) for several state-of-the-art solvers alongside our proposed solvers. Our solvers exhibit significantly faster convergence and attain lower energy configurations compared to existing techniques. The Free Energy Machine (FEM) was evaluated using its official implementation [fem2025] on an NVIDIA Jetson Nano.
@@
Figure S2: Benchmark results for a $10^{4}$ spin SK model. The plot shows Ising energy versus computation time on a logarithmic scale for a range of established and proposed solvers. Our solvers consistently demonstrate advantages in both convergence speed and energy minimization as the problem size increases. FEM results are obtained using the official implementation [fem2025], executed on an NVIDIA Jetson Nano.
@@
Figure S3: Benchmark results for a large-scale sparse Ising instance with $10^{4}$ spins and $1\%$ connectivity. Non-zero couplings $J_{ij}$ are drawn uniformly from $9$-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). Ising energy is plotted against computation time on a logarithmic scale. The results highlight the scalability and efficiency of our solvers under low-connectivity and discrete-weight regimes. All experiments were executed on a NVIDIA Jetson Nano.
@@
Figure S4: Performance comparison on K100000 (a fully connected Ising model with $10^{5}$ spins and binary couplings $J_{ij}\in\{-1,+1\}$ drawn with equal probability). Top: Ising energy evolution over time (log scale). Bottom: Ising energy as a function of iteration count. These results demonstrate the rapid energy descent of our solvers on large and dense instances. All experiments were performed on a single NVIDIA RTX 3090 GPU.
@@
Figure S5: Performance comparison on a fully connected $10^{5}$ spin SK. The Ising energy is plotted against the iteration count, up to $100$ iterations. We used a single NVIDIA RTX 3090 GPU for the computations.
@@
Figure S6: Performance comparison of Ising energy for a $10^{5}$ spin Ising model with $1\%$ connectivity. Couplings $J_{ij}$ are drawn uniformly from signed 99-bit integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). The results are for a single NVIDIA RTX 3090 GPU.
@@
Figure S7: Time evolution of Ising energy for a $10^{6}$ spin fully connected Ising model with structured couplings $J_{ij}=\sin(i\,j+\mathrm{seed})$. The results are for a single NVIDIA V100 GPU.
@@
Figure S8: Time evolution of Ising energy for a $10^{6}$ spin sparsely connected Ising model, where $0.1\%$ of the couplings are nonzero. Each active coupling $J_{ij}$ is sampled uniformly from the set of 99-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). All solvers were run on a single NVIDIA V100 GPU.
@@
Figure S9: Time evolution of Ising energy for a $10^{6}$ spin extremely sparse Ising model with $0.01\%$ connectivity. Each nonzero coupling $J_{ij}$ is drawn uniformly from the set of 99-bit signed integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). All solver experiments were executed on a single NVIDIA V100 GPU.
@@
Figure S10: Benchmarking performance on a $10^{7}$ spin sparse Ising model with $0.001\%$ connectivity. Coupling coefficients $J_{ij}$ are drawn uniformly at random from the set of signed 9-bit integers ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$). The vertical axis represents the Ising energy, while the horizontal axis (on a log scale) denotes computation time in seconds. Each solver was executed using four NVIDIA V100 GPUs.
@@
Figure S11: Benchmarking performance on a $10^{8}$-spin extremely sparse Ising model with connectivity of only $0.00001\%$. Each nonzero coupling $J_{ij}$ is drawn uniformly at random from the set of signed 9-bit integers, $\{-2^{9}+1,\dots,2^{9}-1\}$. The algorithms were run on two NVIDIA H100 GPUs. The Ising energy is plotted as a function of computation time (in minutes).
Figure S12: MAX-CUT on G-set graphs. The mean cut ratio (mean cut normalized by the cut value reported in [maxcut2017multiple]) is shown for G-set graphs (random: $G_{6},G_{7},G_{8}$; toroidal: $G_{11},G_{12},G_{13}$; planar: $G_{18},G_{19},G_{20},G_{21}$). Panels (a), (b), and (c) show the results after 55, $10$, and $100$ iterations. The error bars indicate the maximum and minimum values across $100$ random initializations, respectively. The computations were performed on a NVIDIA RTX 3050 GPU laptop.
Figure S13: Error-bar analysis for medium and large-scale Ising models. The plots show the Ising energy ratio (energy obtained by a solver normalized by the best energy across solvers) after a fixed runtime $T$. For each solver, the mean energy ratio is shown, with error bars indicating the maximum and minimum energy values over $100$ random initializations. (a) $10^{4}$-spin SK model, runtime $T=1$ s; (b) $10^{4}$-spin Ising model with $1\%$ connectivity, runtime $T=1$ s. Couplings $J_{ij}$ are drawn uniformly from signed 9-bit integers ($J_{ij}\in{-2^{9}+1,\dots,2^{9}-1}$); (c) $K100000$ graph, $T=10$ s; (d) $10^{5}$-spin SK model, $T=10$ s; (e) $10^{5}$-spin Ising model with $1\%$ connectivity, runtime $T=10$ s. Couplings $J_{ij}$ are drawn uniformly from signed 9-bit integers ($J_{ij}\in{-2^{9}+1,\dots,2^{9}-1}$). All the computations were performed on a single NVIDIA RTX 3090 GPU.
@@
Figure S14: Battery-powered edge computing setup based on the NVIDIA Jetson Nano module. The system includes a display, keyboard, and mouse interfaced with the Jetson board, powered through a $60,000$ mAh portable power system.
@@
We can represent any $n$-spin Ising model with an external field as an $(n+1)$-spin Ising model without an external field. Consider the Ising model with an external field:
@@
Let $\boldsymbol{s}\in\{-1,1\}^{n}$ and $t\in\{-1,1\}$. From (S1) and (S2),
@@
On the other hand, by the definition of $\boldsymbol{\sigma}^{*}$,
@@
Comparing (S3) and (S4), we obtain
$\mathcal{E}(\boldsymbol{s}^{*})=\mathcal{E}(t_{0}\boldsymbol{s}_{0})=\hat{\mathcal{E}}(\left(\boldsymbol{s}_{0},t_{0}\right))\leqslant\hat{\mathcal{E}}(\left(\boldsymbol{s},1)\right)=\mathcal{E}(\boldsymbol{s})$
Since this holds for any $\boldsymbol{s}\in\{-1,1\}^{n}$, it follows that $\boldsymbol{s}^{*}$ is a ground state of (S1). ∎
@@
We replace the binary vector $\boldsymbol{s}$ with $\boldsymbol{x}\in\mathbb{R}^{n}$ and consider the corresponding relaxed energy
@@
To bias the components of $\boldsymbol{x}$ toward binary values, we introduce the attractor
$\mathcal{A}(\boldsymbol{x})=\frac{\beta}{4}(x_{1}^{4}+\dots+x_{n}^{4})-\frac{\alpha}{2}(x_{1}^{2}+\dots+x_{n}^{2}).$ (S7)
@@
In other words, the attractor penalizes deviations from $\pm\sqrt{\alpha/\beta}$. This property is useful because it forces the relaxed continuous variables toward values that closely mimic the original binary spins, thereby guiding the optimization process toward candidate ground states of the Ising model.
We form the Hamiltonian by combining the Ising energy $\mathcal{E}$ with the attractor $\mathcal{A}$:
$\mathcal{H}=\mathcal{A}+\mathcal{E}.$ (S10)
which forms the basis for our proposed Ising solver. The key observation behind our algorithm is that we can write (S10) as a difference of convex functions:
$\mathcal{H}=f-g,$
where
@@
The function $f$ is convex for all $\beta>0$, and there exists $\alpha_{0}$ such that $g$ is strongly convex for all $\alpha>\alpha_{0}$.
@@
The Hamiltonian $\mathcal{H}$ is coercive for all $\beta>0$, that is, $\mathcal{H}(\boldsymbol{x})\to+\infty$ as $\|\boldsymbol{x}\|_{2}\to\infty$.
@@
This follows from the Weierstrass extreme value theorem [rudin1987real], since $\mathcal{H}$ is continuous and coercive. ∎
@@
$\mathcal{E}(\mathrm{sign}(\boldsymbol{x}^{*}))\leqslant\mathcal{E}(\boldsymbol{s}^{*}).$ (S12)
@@
Since $\boldsymbol{x}^{*}\in\{-\lambda,+\lambda\}^{n}$, we have $\boldsymbol{x}^{*}=\lambda\,\mathrm{sign}(\boldsymbol{x}^{*})$. Setting $\boldsymbol{x}^{*}=\lambda\,\mathrm{sign}(\boldsymbol{x}^{*})$ and $\boldsymbol{x}=\lambda\boldsymbol{s}^{*}$ in (S13) gives
@@
From (S6) and  (S7), it follows that $\mathcal{E}(\lambda\boldsymbol{x})=\lambda^{2}\mathcal{E}(\boldsymbol{x})$ and $\mathcal{A}(\lambda\,\mathrm{sign}(\boldsymbol{x}^{*}))=\mathcal{A}(\lambda\boldsymbol{s}^{*})$. Substituting these into (S14), we get (S12). ∎
@@
In general, the challenging part of convergence analysis is to show that the iterates remain bounded. This is often assumed in the analysis [phan2018]. However, for DOCH, boundedness of the iterates can be established unconditionally.
@@
By Proposition 6, the sequence $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$ is monotone (nonincreasing). Hence, for all $k\geqslant 1$,
@@
As $\mathcal{H}$ is coercive, the sublevel set $\{\boldsymbol{x}:\mathcal{H}(\boldsymbol{x})\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})\}$ is bounded. However, by (S21), the sequence $\{\boldsymbol{x}^{(k)}\}$ is contained in this sublevel set and is hence bounded.
@@
Since $\mathcal{H}(\boldsymbol{x}^{(k)})$ is bounded, there exists $C>0$ such that $\mathcal{H}(\boldsymbol{x}^{(k)})\geqslant C$ for all $k\geqslant 1$. Hence,
$\frac{\mu}{2}\sum_{k=0}^{N}\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k+1)}\|_{2}^{2}\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})-\mathcal{H}(\boldsymbol{x}^{(N+1)})\leqslant\mathcal{H}(\boldsymbol{x}^{(0)})-C<+\infty.$
Letting $N\to\infty$ gives
$\frac{\mu}{2}\sum_{k=0}^{\infty}\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{(k+1)}\|_{2}^{2}<+\infty.$ (S22)
Consequently, we obtain (S20). ∎
Note that (S22) by itself does not guarantee that the sequence $\{\boldsymbol{x}^{(k)}\}$ converges. To prove convergence, we must show that $\{\boldsymbol{x}^{(k)}\}$ is a Cauchy sequence [rudin1987real]. This will be done in Theorem 2, using the fact that the Hamiltonian $\mathcal{H}$ is real-analytic (being a polynomial) and therefore satisfies the Łojasiewicz gradient inequality [KL1963]. Based on the analysis so far, we can already state the following result.
@@
The iterates $\{\boldsymbol{x}^{(k)}\}$ are bounded by Proposition 7, and hence, by the Bolzano-Weierstrass theorem [rudin1987real], have at least one limit point, say, $\boldsymbol{x}^{*}\in\mathbb{R}^{n}$. In particular, we can extract a subsequence $\{\boldsymbol{x}^{(k_{i})}\}$ such that
@@
Letting $i\to\infty$ on both sides using (S22) gives
@@
We now use the classical Łojasiewicz gradient inequality [KL1963] to show that the DOCH iterates actually converge, that is, they admit a unique limit point. For this, we need the following result.
@@
Letting $\sigma=\lambda_{\max}(\mathbf{J}+\alpha\mathbf{I})>0$, we get (S27). ∎
The following is the main convergence result. The proof follows the analysis in [absil2005, bolte_kl_2013].
@@
From Proposition 6, we know that $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$ is non-increasing. Moreover, since $\boldsymbol{x}^{*}$ is a limit point of $\boldsymbol{x}^{(k)}$ and $\mathcal{H}$ is continuous,
@@
Since $\mathcal{H}$ is a polynomial, it satisfies the Łojasiewicz gradient inequality [KL1963, absil2005] in some neighbourhood of the critical point $\boldsymbol{x}^{*}$. More specifically, there exist $\theta\in[0,1),c>0$ and $r>0$, such that for all $\|\boldsymbol{x}-\boldsymbol{x}^{*}\|<r$,
@@
We can conveniently reformulate this in terms of the function
$\varphi(t)=\frac{c}{1-\theta}\,t^{1-\theta}.$ (S30)
We can verify that (S30) is concave and non-decreasing on $0,\infty)$, and $\varphi^{\prime}(t)=ct^{-\theta}$ for any $t>0$. Therefore, we can write (S29) as
@@
We know from (S20) and (S28) that $\delta_{k}\to 0$ and $\Delta_{k}\to 0$. Since $\varphi$ is non-decreasing and satisfies $\varphi(t)\to 0$ as $t\to 0$, it follows that $\varphi(\Delta_{k})\to 0$. Moreover, since $\boldsymbol{x}^{*}$ is a limit point of $\{\boldsymbol{x}^{(k)}\}$, there exists $k_{0}\geqslant 1$ such that
@@
On the other hand, by (S16), we have
@@
Combining this with (S32) gives
@@
From (S32) and (S36), we have
@@
Thus $\boldsymbol{x}^{k_{0}+1}$ is in the neighbourhood of $\boldsymbol{x}^{*}$ stipulated in (S29). In fact, we claim that the entire tail $\{\boldsymbol{x}^{k}\},k>k_{0}$, belongs to this neighbourhood. This can be shown using contradiction. Indded, suppose $k_{1}$ is the smallest $k>k_{0}$ such that $\|\boldsymbol{x}^{(k_{1})}-\boldsymbol{x}^{*}\|_{2}\geqslant r$. This means that the points $\boldsymbol{x}_{k_{0}},\ldots,\boldsymbol{x}_{k_{1}-1}$ are with the neighbourhood stipulated in (S29). Applying (S35) repeatedly, we have
$\delta_{k+1}\leqslant\frac{2\sigma}{\mu}\bigl(\varphi(\Delta_{k})-\varphi(\Delta_{k+1})\bigr)+(\delta_{k}-\delta_{k+1}),\qquad(k_{0}\leqslant k<k_{1}).$
Summing both sides, we obtain